OpenStudy (anonymous):

Find the coordinate(s) of the points on the curve x^3 − 3x^2 where the tangent: a) is parallel to the x-axis b) is parallel to the line y = −3x + 2.

OpenStudy (anonymous):

(a) parallel to the x axis means the same gradient.. so gradient is defined as (y2-y1)/(x2-x1), ur y value would be the same as its just a horizontal line, so ur gradient have to be 0

OpenStudy (anonymous):

differentiate x^3 -3x^2.. u get 3x^2 - 6x = 0 solve for x.. 3x^2= 6x x^2=2 x= + root 2 n negative root 2

OpenStudy (anonymous):

once u get ur x values then sub in to ur original equation to get ur y coordinate

OpenStudy (anonymous):

but the answers in the back of the book don't have the x values you stated

OpenStudy (anonymous):

for (b), differentiate the line y = -3x + 2, u get -3 then 3x^2 - 6x = -3 solve for x... 3(x^2 -2x +1 )=0 3(x-1)(x-1)=0 x has to be equal to negative 1.. hten sub in to ur equation and get ur y coordinate

OpenStudy (anonymous):

whats the answer at the back?

OpenStudy (anonymous):

(0,0) and (2,-4)

OpenStudy (anonymous):

3x(x-2) = 0

OpenStudy (anonymous):

sorry i calculated wrongly

OpenStudy (anonymous):

when u reached 3x^2 - 6x = 0, u factorize the common factor out 3x(x-2)=0 so u get x =2 and x = 0

OpenStudy (anonymous):

but u get what i am doing ?

OpenStudy (anonymous):

some, well I got the x=0 or x=2 after factorizing the common factor, just as u did.

OpenStudy (anonymous):

but may I ask, why do we have to solve for x?

OpenStudy (anonymous):

u have to solve for x cause u want ur x coordinate o.O? then u can sub ur x coordinate back to the equation x^3 - 3x^2 to get ur y coordinate

OpenStudy (anonymous):

yeh and what is the purpose of it then*? I know it's to find the x and y coordinate and the point ultimately.

OpenStudy (anonymous):

dnt really get what u mean by whats the purpose.. that point tells u that the gradient is 0

OpenStudy (anonymous):

don't worry.

OpenStudy (anonymous):

sry :(

OpenStudy (anonymous):

it's fine, it's just me, cause I get quite confused sometimes and I like to or at least try to understand things properly.

OpenStudy (anonymous):

in a logical order* if that makes sense.

OpenStudy (anonymous):

a) parallel to the x-axis = gradient equal to zero gradient = zero when the function is at a maximum or a minimum to find the max/min of the function, differentiate the function so dy/dx= 3x^2 - 6x make it equals to zero 0= 3x^2-6x 0=x(3x-6) so x= 0 or 3x-6=0 for that we get x=0 or x=2 so sub the x value into the equation y=x^3-3x^2 u get y=0 when x=0 coordinate = (0,0) y=-4 when x=2 coordinate = (2,-4)

OpenStudy (anonymous):

b) parallel to the line y=-3x+2 gradient must be = -3 to be parallel dy/dx is the gradient function of the function y so to find the coordinate: dy/dx= 3x^2-6x m(gradient)=-3 so -3=3x^2-6x (gradient of line equals to the gradient of the tangent of that point) solve that we get x=1 ( the x value for that coordinate) so sub x=1 into the function y=x^3-6x we get y=-2 so coordinate = (1,-2)